# CORT: Class-Oriented Real-time Tracking for Embedded Systems

### v1

@@ Abstract
The ever-increasing use of artificial intelligence in autonomous systems has significantly contributed to advance the research on multi-object tracking, adopted in several real-time applications (e.g., autonomous driving, surveillance drones, robotics) to localize and follow the trajectory of multiple objects moving in front of a camera. Current tracking algorithms can be divided into two main categories: some approaches introduce complex heuristics and re-identification models to improve the tracking accuracy and reduce the number of identification switches, without particular attention to the timing performance, whereas other approaches are aimed at reducing response times by removing the re-identification phase, thus penalizing the tracking accuracy. This work proposes a new approach to multi-class object tracking that allows achieving smaller and more predictable execution times, without penalizing the tracking performance. The idea is to reduce the problem of matching predictions with detections into smaller sub-problems by splitting the Hungarian matrix by class and invoking the second re-identification stage only when strictly necessary for a smaller number of elements. The proposed solution was evaluated in complex urban scenarios with several objects of different types (as cars, trucks, bikes, and pedestrians), showing the effectiveness of the multi-class approach with respect to state of the art trackers.
Index Terms:
Real-time tracking, multiple-object tracking, autonomous driving, embedded systems, real-time computing.
@@
I Introduction
Object tracking is a computer vision task for locating objects in a video sequence and generating their trajectories, associating each object not only with a bounding box, but also with a unique identifier. It is used in many application fields, including surveillance, autonomous driving, robotics, video analysis, and augmented reality. Differently from object detectors, which only analyze single frames, trackers process video sequences to reconstruct the trajectories of the moving objects (tracklets), which are used to re-identify objects in consecutive video frames, even in the presence of occlusions or crossing trajectories.
@@ I Introduction
Tracking by detection has rapidly spread into various application domains, including embedded systems, showing its potential in advanced driving assistance systems (ADAS) [12] and unmanned aerial vehicles (UAV) [13]. However, the integration of this technology into embedded devices faces a major issue due to the limited computational resources available in such systems. Most of the state-of-the-art approaches [5, 14, 15] focused on maximizing the tracking performance on common benchmarks, like the MOT-Challenge [16], introducing complex heuristics and re-identification models to improve the tracking accuracy and reduce the number of identification switches, without particular attention to the timing performance. Other approaches [17, 18] reduced the response time of the tracking pipeline by removing the neural network for the re-identification phase, thus penalizing the tracking accuracy.
@@
  • It proposes a parallel formulation of the matching problem by splitting the Hungarian matrix into smaller ones, using the object class as discriminant. As illustrated in Figure 1, this allows reducing the execution time of the overall matching phase to the one required to solve the largest sub-matrix. Due to the cubic complexity of the Hungarian matching algorithm, the proposed approach is particularly effective in the presence of objects of different classes.
@@
Paper structure The rest of the paper is organized as follows: Section II discusses the state of the art solutions related to this work; Section III presents the proposed approach; Section IV reports some experimental results aimed at showing the benefits of the proposed solution; and, finally, Section V states the conclusions and possible future developments.
II Related work
@@ II-A Joint trackers
Another well-known problem in detection-based trackers is the bias between the detection and the re-identification tasks. FairMOT [24] improved the performance of the tracking pipeline by implementing these two tasks within the same model using parallel branches. This approach solves the natural imbalance between the two tasks, making them ”fair” to each other. It also uses CenterNet [22] as a baseline, while introducing significant improvements in the model. For instance, the use of deformable convolutions [25] helped enhancing the receptive fields in the detection step, while the adoption of the Deep Layer Aggregation (DLA) [26], an enhanced version of the Feature Pyramid Networks (FPN) [27], improved the re-identification task significantly reducing the number of ID switches.
Transformers [28] were also used to learn complex tasks combining different steps of the object tracking pipeline. For instance, TransTrack [29] and Trackformer [30] are great examples of the use of such a technology. Both are based on the Detection Transformers (DETR) [31], which is one of the most prominent approaches for object detection. In particular, TransTrack [29] adds another transformer decoder to the DETR [31] architecture and introduces a matching head for producing tracking results. Trackformer [30] is an improvement of TransTrack [29] that uses just one transformer decoder, which directly produces the tracking queries. The set of queries obtained at time $t-1$ is then used in the current frame to match existing tracks or create new ones. However, transformers are complex models that require substantial computational resources to run in real-time, often not available in embedded devices. As also shown in the TransTrack paper, varying the number of decoders between 1 and 6, decreases the speed from 15 fps to 10 fps on a single V100 GPU, while the Trackformer executes at 7.4 fps.
II-B Separate trackers
@@
To address this problem, the same authors proposed Deep-OC-SORT [15], which introduces an adaptive appearance similarity-based association model upon OC-SORT [18], improving the tracking performance on the DanceTrack [34] benchmark by 6 HOTA [35] points. It also resulted to be the best model, according to the HOTA metrics, on both MOT17 [36] and MOT20 [16] benchmarks, clearly showing the importance of having a strong re-identification model. To date, state of the art re-identificator techniques are all based on deep neural networks, which by far outperform handcrafted or pure mathematical-based strategies. Their main drawback, however, is their computational cost.
This work proposes a new way to formulate the MOT problem more efficiently by splitting the Hungarian algorithm into multiple class-based matching tasks, which can then be parallelized, thus significantly reducing the overall computation time of the algorithm. Re-identification is used in the proposed tracking pipeline, due to its proven effectiveness, but the number of its invocations is reduced to save computation time thanks to a two-stage cascade matching. It worth observing that the proposed solution is general and can also be adopted to other state-of-the-art trackers to improve their inference time without affecting their accuracy.
III Proposed approach
As mentioned above, the proposed solution (named CORT) consists of splitting the Hungarian algorithm into a number of smaller class-based matching tasks, which can be parallelized to reduce the overall computation time of the algorithm. In order to better understand the proposed idea, the following section briefly recalls the Hungarian algorithm used in most separate trackers to optimally match bounding boxes detected at the current frame with bounding boxes predicted in the previous frame.
III-A The Hungarian match
@@ III-B Concurrent MOT formulation
Consider, for instance, the typical situation in which each prediction has a corresponding detection to be associated with ($p=d$). In this case, all the objects are successfully tracked and there are no tracks to be deleted or created. Figure 2 shows how the data are organized in the traditional approach (left) and in the proposed solution (right).
@@
In the example illustrated in Figure 2, colors are associated with classes and it is clear that, when the objects in the scene belong to different classes, the number of values to be computed in the standard formulation is higher than the one required for the proposed formulation. Such a difference tends to increase with the number of objects to be tracked, hence the advantage of the class-based formulation tends to become more significant. A quantitative analysis of the resources required by the two approaches is presented in the next section.
Now, consider a situation in which the number of detections is higher than the number of predictions ($d>p$), because a new object entered the scene and requires the creation of a new tracklet. This is a very common situation that frequently occurs in a MOT task. For instance, assume that there are four predictions from the previous frame and five detections in the current frame, and assume that all the pre-existing tracks are successfully associated with one detection. This example is illustrated in Figure 3.
@@
Adding a new element in the standard cost matrix involves creating a new row ($P_{5}$), filling all its elements with the dummy cost $V$, and then computing the similarity score between the new prediction ($D_{5}$) and all the pre-existing predictions $(P_{1},\ldots,P_{4})$.
@@
III-C Remarks on resource saving
@@
In the case considered in Figure 2, the solution of the standard Hungarian match has a time complexity proportional to $2166^{3}=$ 216 steps, while the proposed class-based solution only requires $243*2^{3}=$ 24 steps, which reduce to 8 steps, if the three matching sub-problems can run in parallel on three different cores.
@@
III-D Proposed tracking pipeline
@@
Figure 4 shows the workflow of the two-stage cascade matching implemented in a generic thread at each iteration of the tracking loop. At the first stage, each cost matrix is populated only using the IoU association metric, where the cost computed for a generic (prediction, detection) pair $(P_{i},D_{j})$ is given by
@@
where $f_{ik}$ is the $k$-th feature array in the history for track $i$, while $f_{j}$ is the feature array extracted from the current detection $j$. Resolving associations in the second stage based not only on the object position, but also using an appearance model, is crucial to increase the tracking performance and allows the tracker to recover from mid/long term occlusions.
IV Experiments
@@
IV-A Evaluating time and memory scalability
@@
Figure 1 shows the ideal scenario for CORT generated in the experiment, where six objects are present in the scene, equally distributed among three classes (car, bus, motorbike). In this case, the traditional Hungarian match formulation requires the creation of a $6\times 6$ cost matrix for a total of 36 entries representing the association scores between predictions and detections. Using the class-based approach, the problem requires the use of three $2\times 2$ cost matrices, for a total of 12 entries (4 + 4 + 4).
Table I reports the execution times required to populate the Hungarian matrix (or matrices, for multi-class approach) and solve the assignment problem(s). Each number reported in the table is the average over 50 executions. The two approaches have been executed in seven different scenarios, each identified as ($N_{1}$, $N_{2}$, $N_{3}$), where $N_{i}$ denotes the number of objects of class $C_{i}$ present in the scene ($C_{1}$ = ”car”, $C_{2}$ = ”bus”, $C_{3}$ = ”motorbike”).
It is worth observing that, in the best-case scenario (2, 2, 2), the time reduction reported for the proposed method is caused not only by the parallel execution of the threads in the corresponding cores of the processing platform, but also by the reduced size of the matrices, as explained in Section III. Although one could expect similar execution times for the three involved classes, this is not necessarily true. In fact, in CORT, the execution time depends both on the number of objects in the class and how many of them are forwarded to the second matching stage.
@@
TABLE I: Execution times in milliseconds for processing the Hungarian match in the proposed multi-class parallel approach (CORT) and in the monolithic DeepSORT solution.
Under more general scenarios, when the number of objects in each class is unbalanced, the execution time of the multi-class approach is given by the largest class matching problem. Figure 5 shows a frame of the CARLA video sequence where there are three cars, two buses, and one motorbike.
@@
The results of this test are also reported in Table I at the row labeled as (3, 2, 1). For the nature of the parallel solution, we can also say that increasing the number of objects in the other classes does not affect the timing performance of the algorithm significantly, as long as their number does not exceed the number of elements in the majority class and the new elements do not require intensive use of the second stage of the pipeline described in Section III. In contrast, in the monolithic approach, increasing the number of elements, regardless of their class, always increases the overall computation time of the algorithm. This is confirmed by the results reported in Table I; in fact, the execution time measured for the proposed multi-class approach for the case (3, 3, 3) is comparable with the time achieved for (3, 2, 1), whereas for the monolithic approach, the time for the case (3, 3, 3) results to be larger than the one for (3, 2, 1). This trend becomes even more evident in the (4, 4, 4) scenario, which includes 12 objects in the scene. Other two tests were carried out using different class distributions, as (6, 0, 0) and (0, 0, 6). Note that these scenarios represent the worst-case situation for the proposed approach, since all the elements belong to the same class. In this case, the resulting execution time of the multi-class method is reduced only by the different implementations of the cascade matching. In fact, for the (6, 0, 0) scenario, the proposed solution can solve the association problem using just the IoU matching for the majority of the frames, saving the time required by the second stage operations. This happens because cars are easier to detect with respect to motorbikes, leading to high confidence predictions from the Kalman filter. Vice versa, motorbikes, being smaller than cars and buses, are more likely to fail the first stage matching based on bounding box localization, so they are passed to the second stage, based on the appearance. This is why the execution times recorded for the (0, 0, 6) scenario are larger than those found for the (6, 0, 0) scenario, even though the number of objects is exactly the same. Finally, note that from the DeepSORT perspective, the execution times of the (6, 0, 0) and the (0, 0, 6) scenarios are comparable, since all the objects undergo the same computation, independently of their class.
IV-B Comparison with BYTEtrack
This experiment has been carried out to compare the execution time and the accuracy of CORT with the one of the best state of the art algorithm for real-time applications, that is BYTEtrack.
The main difference between the two tracking algorithms is the way they formulate the Hungarian match. In fact, BYTEtrack uses a decision threshold to split the detections into high and low confidence ones, and it never uses a neural network to perform a feature-based re-identification step. Differently, the proposed solution first divides the problem by classes and then performs the two-stage cascade matching explained in Section III. It is worth noticing that the choice of not using a neural network to get an appearance model exposes BYTEtrack to a higher number of id-switches when objects overlap to each other or in the presence of long-term occlusions.
Consider, for example, the scenario represented in the three frames reported in Figure 6.

### v2

@@ Abstract
The ever-increasing use of artificial intelligence in visual perception tasks for autonomous systems has significantly contributed to advance the research on multi-object tracking, which is a function required in several real-time applications (e.g., autonomous driving, surveillance drones, robotics) to localize and follow the trajectory of multiple objects moving in front of a camera. Most of current tracking algorithms introduce complex heuristics and re-identification models to improve the tracking accuracy and reduce the number of identification switches, without particular attention to the timing performance, whereas other approaches are aimed at reducing response times by removing the re-identification phase, thus penalizing the tracking accuracy. This work proposes a new approach to multi-class object tracking that allows achieving smaller and more predictable execution times with respect to traditional approaches, without penalizing the tracking performance. The idea is to divide the problem of matching predictions with detections into a number of smaller sub-problems by splitting the Hungarian association matrix by class and invoking the second re-identification stage only when strictly necessary, thus applying it to a smaller number of elements. Splitting the matching problem into a number of smaller sub-problems also allows parallelizing the Hungarian algorithm, further reducing the execution time in multi-core processing platforms. The proposed solution was evaluated in complex urban scenarios with different types of objects (as cars, buses, and motorbikes), and different number of instances, showing the effectiveness of the multi-class approach in reducing execution times without penalizing performance, with respect to state of the art trackers.
{IEEEImpStatement}
The proposed class-based solution helps containing the execution time of the tracking pipeline without penalizing the performance. In fact, the association problem is divided into a number of smaller sub-problems (equal to the number of object classes present in the scene) that are simpler to solve and can also be executed in parallel on a multi-core architecture. This allows the entire tracking pipeline to be executed at higher rates also on small embedded platforms, thus enabling object tracking to be used in a variety of applications for unmanned aerial and terrestrial vehicles. The proposed approach is general and can be adopted in any tracking-by-detection algorithm.
{IEEEkeywords}
Artificial intelligence, Object tracking, Performance optimization, Real-time tracking, Tracking algorithm, Tracking by detection.
@@
1 Introduction
[…]
Object tracking is a computer vision task for locating objects in a video sequence and generating their trajectories, associating each object not only with a bounding box, but also with a unique identifier. It is used in many application fields, including surveillance, autonomous driving, robotics, video analysis, and augmented reality. In autonomous driving, object tracking is crucial for detecting and monitoring pedestrians, vehicles, and other dynamic objects in real-time, ensuring safe navigation by providing reliable, continuous object information in complex scenes. In surveillance, tracking systems enable anomaly detection, which is essential for identifying unusual behaviors and enhancing security in crowded or high-traffic environments.
@@ 1 Introduction
Tracking by detection has rapidly spread into various application domains, including embedded systems, showing its potential in advanced driving assistance systems (ADAS)[12] and unmanned aerial vehicles (UAV)[13]. However, the integration of this technology into embedded devices faces a major issue due to the limited computational resources available in such systems. Most of the state-of-the-art approaches [5, 14, 15] focused on maximizing the tracking performance on common benchmarks, like the MOT-Challenge [16], introducing complex heuristics and re-identification models to improve the tracking accuracy and reduce the number of identification switches, without particular attention to the timing performance. Other approaches [17, 18] reduced the response time of the tracking pipeline by removing the neural network for the re-identification phase, thus penalizing the tracking accuracy. In some applications, as autonomous driving and surveillance, increasing the efficiency of multi-class tracking is essential to handle diverse object types and meet the real-time requirements necessary to manage dynamic and crowded environments effectively.
@@
  • It proposes a parallel formulation of the matching problem by splitting the Hungarian matrix into smaller ones, using the object class as discriminant. As illustrated in Fig. 1, this allows reducing the execution time of the overall matching phase to the one required to solve the largest sub-matrix. Due to the cubic complexity of the Hungarian matching algorithm, the proposed approach is particularly effective in the presence of objects of different classes.
@@
Paper structure The rest of the paper is organized as follows: Section 2 discusses the state of the art solutions related to this work; Section 3 presents the proposed approach; Section 4 reports some experimental results aimed at showing the benefits of the proposed solution; and, finally, Section 5 states the conclusions and possible future developments.
2 Background
@@ 2.1 Joint trackers
Another well-known problem in detection-based trackers is the bias between the detection and the re-identification tasks. FairMOT [24] improved the performance of the tracking pipeline by implementing these two tasks within the same model using parallel branches. This approach solves the natural imbalance between the two tasks, making them ”fair” to each other. It also uses CenterNet [22] as a baseline, while introducing significant improvements in the model. For instance, the use of deformable convolutions [25] helped enhancing the receptive fields in the detection step, while the adoption of the Deep Layer Aggregation (DLA) [26], an enhanced version of the Feature Pyramid Networks (FPN) [27], improved the re-identification task by aggregating multi-scale features for consistent embeddings across frames, even under size variation or partial occlusion. This unified architecture reduces ID switches by handling detection and re-identification jointly, making tracking more robust to rapid changes in object appearance.
Transformers [28] were also used to learn complex tasks combining different steps of the object tracking pipeline. For instance, TransTrack [29] and Trackformer [30] are great examples of the use of such a technology. Both are based on the Detection Transformers (DETR) [31], which is one of the most prominent approaches for object detection.
In particular, TransTrack [29] adds another transformer decoder to the DETR [31] architecture and introduces a matching head for producing tracking results. Such a dual-decoder design allows associating object detections across frames by learning spatial and temporal dependencies, thus enhancing tracking accuracy; however, it requires substantial computational resources, limiting its real-time applicability on embedded systems. Trackformer [30] uses just one complex transformer decoder to directly produce the tracking queries. The set of queries obtained at the previous frame is then used in the current frame to match existing tracks or create new ones, simplifying the process but increasing the decoder complexity. Although Trackformer increases tracking accuracy while preserving the transformer-based advantage of modeling long-term dependencies, it still requires a high computational power, often not available in small embedded platforms. In particular, as reported in the corresponding papers, the speed of TransTrack on a single V100 GPU varies from 10 fps to 15 fps when reducing the number of decoders from 6 to 1, while the Trackformer executes at 7.4 fps.
2.2 Separate trackers
@@
To address this problem, the same authors proposed Deep-OC-SORT [15], which introduces an adaptive appearance similarity-based association model upon OC-SORT [18], improving the tracking performance on the DanceTrack [34] benchmark by 6 HOTA [35] points. It also resulted to be the best model, according to the HOTA metrics, on both MOT17 [36] and MOT20 [16] benchmarks, clearly showing the importance of having a strong re-identification model.
For this reason, many modern multi-object trackers leverage complex neural networks for re-identification, such as those based on Siamese network architectures [37]. While these methods greatly improve tracking performance by providing more robust and accurate identity preservation, they impose significant limitations on their use in embedded systems and real-time applications due to the high computational demands and latency introduced by such sophisticated models. Because of that, various alternative approaches have been explored to reduce the computational load. For instance, QDTrack [38] employs a quasi-dense strategy by sampling hundreds of regions for contrastive learning. While this method has proven effective, it achieves a processing speed of 20.3 fps, which may not be sufficient for certain real-world applications, particularly those involving fast-moving objects, such as drones or autonomous vehicles.
A different strategy is adopted by Fast re-obj [39], which combines segmentation and embedding modules to obtain fast and reliable association features. However, although the approach is effective for real-time systems, the limitations imposed by the rigid scenes, makes it not suitable for dynamic environments. To solve the problems highlighted above, this paper proposes a new approach that is able to reduce the execution time of the algorithm for matching detections and predictions in the tracking pipeline, under complex scenarios including several moving objects belonging to different categories. Objects can overlap or be temporarily occluded by other objects. The proposed approach splits the Hungarian algorithm into multiple class-based matching tasks, which can then be parallelized to significantly reduce the overall computation time. To balance accuracy and efficiency, our tracking pipeline incorporates re-identification; however, to contain the overall computational load, the number of re-identification invocations is controlled through a two-stage cascade matching process. It is worth noting that the proposed solution is general and can be adopted by other state-of-the-art trackers to improve inference time without sacrificing accuracy.
3 Methodology
As mentioned above, the problem addressed in this work is the high computational cost of traditional multi-object trackers, which limits their use in real-time applications on resource-constrained embedded systems. In particular, the monolithic formulation of the Hungarian algorithm, required to match detections with predictions, scales poorly as the number of objects in the scene increases, due to the cubic complexity of the algorithm and the time required to create and delete new tracks. To address this issue, the proposed solution (named CORT) reformulates the Hungarian algorithm into a number of smaller class-based matching tasks, which can be parallelized to reduce the overall computation time of the algorithm. In order to better understand the proposed idea, the following section briefly recalls the Hungarian algorithm used in most separate trackers to optimally match bounding boxes detected at the current frame with bounding boxes predicted in the previous frame.
3.1 The Hungarian match
@@ 3.2 Concurrent MOT formulation
Consider, for instance, the typical situation in which each prediction has a corresponding detection to be associated with ($p=d$). In this case, all the objects are successfully tracked and there are no tracks to be deleted or created. Fig. 2 shows how the data are organized in the traditional approach (left) and in the proposed solution (right).
@@
In the example illustrated in Fig. 2, colors are associated with classes and it is clear that, when the objects in the scene belong to different classes, the number of values to be computed in the standard formulation is higher than the one required for the proposed formulation. Such a difference tends to increase with the number of objects to be tracked, hence the advantage of the class-based formulation tends to become more significant. A quantitative analysis of the resources required by the two approaches is presented in the next section.
Now, consider a situation in which the number of detections is higher than the number of predictions ($d>p$), because a new object entered the scene and requires the creation of a new tracklet. This is a very common situation that frequently occurs in a MOT task. For instance, assume that there are four predictions from the previous frame and five detections in the current frame, and assume that predictions $P_{1}$ and $P_{4}$ are matched in the first stage, while $P_{2}$ and $P_{3}$ are matched in the second stage of the tracking pipeline. This example is illustrated in Fig. 3.
@@
Adding a new element in the standard cost matrix involves creating a new row ($P_{5}$), filling all its elements with the dummy cost $V$, and then computing the similarity score between the new detection ($D_{5}$) and all the pre-existing predictions $(P_{1},\ldots,P_{4})$.
@@
3.3 Remarks on resource saving
@@
In the case considered in Fig. 2, the solution of the standard Hungarian match has a time complexity proportional to $2166^{3}=$ 216 steps, while the proposed class-based solution only requires $243*2^{3}=$ 24 steps, which reduce to 8 steps, if the three matching sub-problems can run in parallel on three different cores.
@@
3.4 Proposed tracking pipeline
@@
Fig. 4 illustrates the workflow adopted in CORT and the interactions among the major components of the tracking system. In particular, the ”Matching logic” block provides a high-level overview of the two-stage cascade matching: the first stage attempts to match detections and predictions based on the IoU, while at the second stage matching is performed based on appearance vectors, measured using cosine similarity. To introduce a tolerance when creating and deleting a tracklet, two patience values, in_patience and out_patience, are defined by the user. Hence, two counters, in_counter and out_counter are associated with each track. When a new object appears (i.e., there is a detection that does not match with any prediction) it is assigned a new “tentative” track, which becomes “effective” only if the object is detected and matched with a prediction for a given number of consecutive frames (in_patience). Referring to Fig. 4, the Initialization logic block creates a tentative track and adds it to the tracks array. Then, at each iteration, the Track State Manager checks whether the track is active or meets the conditions required for activation. A track becomes effective if in_counter $=$ in_patience. A tentative track that fails in the first matching stage (i.e., it does not satisfy the in_patience criterion) is deleted. Similarly, if an object associated with an effective track is not detected for a subsequent number of frames (out_patience) the corresponding track is deleted. In Fig. 4, this flow is represented by the arrow that, after the second matching stage, connects the Unmatched predictions with the Dead reckoning and Elimination logic. In this block, if the out_counter reaches the out_patience value, the track is deleted, otherwise it is maintained in the Tracks array. At the first stage, each cost matrix is populated only using the IoU association metric, where the cost computed for a generic (prediction, detection) pair $(P_{i},D_{j})$ is given by
@@
where $f_{ik}$ is the $k$-th feature array in the history for track $i$, while $f_{j}$ is the feature array extracted from the current detection $j$. Resolving associations in the second stage based not only on the object position, but also using an appearance model, is crucial to increase the tracking performance and allows the tracker to recover from mid/long term occlusions avoiding erroneous track deletion and re-initialization.
4 EXPERIMENTS
@@
4.1 Evaluating time and memory scalability
@@
Fig. 1 shows the ideal scenario for CORT generated in the experiment, where six objects are present in the scene, equally distributed among three classes (car, bus, motorbike). In this case, the traditional Hungarian match formulation requires the creation of a $6\times 6$ cost matrix for a total of 36 entries representing the association scores between predictions and detections. Using the class-based approach, the problem requires the use of three $2\times 2$ cost matrices, for a total of 12 entries (4 + 4 + 4).
Table 1 reports the execution times required to populate the Hungarian matrix (or matrices, for multi-class approach) and solve the assignment problem(s). Each number reported in the table is the average over 50 executions. The two approaches have been executed in seven different scenarios, each identified as ($N_{1}$, $N_{2}$, $N_{3}$), where $N_{i}$ denotes the number of objects of class $C_{i}$ present in the scene ($C_{1}$ = ”car”, $C_{2}$ = ”bus”, $C_{3}$ = ”motorbike”).
It is worth observing that, in the best-case scenario (2, 2, 2), the time reduction reported for the proposed method is caused not only by the parallel execution of the threads in the corresponding cores of the processing platform, but also by the reduced size of the matrices, as explained in Section 3. Although one could expect similar execution times for the three involved classes, this is not necessarily true. In fact, in CORT, the execution time depends both on the number of objects in the class and how many of them are forwarded to the second matching stage.
@@
Table 1: Execution times in milliseconds for processing the Hungarian match in the proposed multi-class parallel approach (CORT) and in the monolithic DeepSORT solution.
Under more general scenarios, when the number of objects in each class is unbalanced, the execution time of the multi-class approach is given by the largest class matching problem. Fig. 5 shows a frame of the CARLA video sequence where there are three cars, two buses, and one motorbike.
@@
The results of this test are also reported in Table 1 at the row labeled as (3, 2, 1). For the nature of the parallel solution, we can also say that increasing the number of objects in the other classes does not affect the timing performance of the algorithm significantly, as long as their number does not exceed the number of elements in the majority class and the new elements do not require intensive use of the second stage of the pipeline described in Section 3. In contrast, in the monolithic approach, increasing the number of elements, regardless of their class, always increases the overall computation time of the algorithm. This is confirmed by the results reported in Table 1; in fact, the execution time measured for the proposed multi-class approach for the case (3, 3, 3) is comparable with the time achieved for (3, 2, 1), whereas for the monolithic approach, the time for the case (3, 3, 3) results to be larger than the one for (3, 2, 1). This trend becomes even more evident in the (4, 4, 4) scenario, which includes 12 objects in the scene. Other two tests were carried out using different class distributions, as (6, 0, 0) and (0, 0, 6). Note that these scenarios represent the worst-case situation for the proposed approach, since all the elements belong to the same class. In this case, the resulting execution time of the multi-class method is reduced only by the different implementations of the cascade matching. In fact, for the (6, 0, 0) scenario, the proposed solution can solve the association problem using just the IoU matching for the majority of the frames, saving the time required by the second stage operations. This happens because cars are easier to detect with respect to motorbikes, leading to high confidence predictions from the Kalman filter. Vice versa, motorbikes, being smaller than cars and buses, are more likely to fail the first stage matching based on bounding box localization, so they are passed to the second stage, based on the appearance. This is why the execution times recorded for the (0, 0, 6) scenario are larger than those found for the (6, 0, 0) scenario, even though the number of objects is exactly the same. Finally, note that from the DeepSORT perspective, the execution times of the (6, 0, 0) and the (0, 0, 6) scenarios are comparable, since all the objects undergo the same computation, independently of their class.
4.2 Comparison with BYTEtrack
This experiment has been carried out to compare the execution time and the accuracy of CORT with the one of the best state of the art algorithm for real-time applications, that is BYTEtrack. The key distinction between CORT and BYTEtrack lies in how they handle the association step. BYTEtrack, as all other tracking models, employs a monolithic approach for solving the Hungarian matching, where all detections, regardless of their class, are processed in a single assignment step. It also exclusively relies on an IoU-based association strategy, splitting detections into high and low confidence groups, without incorporating a feature-based re-identification mechanism. In contrast, CORT introduces a class-based decomposition of the Hungarian matching process, allowing each class-specific assignment problem to be solved independently and in parallel. Furthermore, CORT refines the association process through a two-stage cascade matching strategy explained in Section 3, where a lightweight CIoU matching is performed first, followed by a feature-based re-identification step, only applied to ambiguous cases. It is worth noticing that the choice of not using a neural network to get an appearance model exposes BYTEtrack to a higher number of id-switches when objects overlap to each other or in the presence of long-term occlusions.
Consider, for example, the scenario represented in the three frames reported in Fig. 6.